### v1

# GenUP: Generative User Profilers as In-Context Learners for Next POI Recommender Systems

Abstract.
Traditional POI recommendation systems often lack transparency, interpretability, and scrutability due to their reliance on dense vector-based user embeddings. Furthermore, the cold-start problem—where systems have insufficient data for new users—limits their ability to generate accurate recommendations. Existing methods often address this by leveraging similar trajectories from other users, but this approach can be computationally expensive and increases the context length for LLM-based methods, making them difficult to scale. To address these limitations, we propose a method that generates natural language (NL) user profiles from large-scale, location-based social network (LBSN) check-ins, utilizing robust personality assessments and behavioral theories. These NL profiles capture user preferences, routines, and behaviors, improving POI prediction accuracy while offering enhanced transparency. By incorporating NL profiles as system prompts to LLMs, our approach reduces reliance on extensive historical data, while remaining flexible, easily updated, and computationally efficient. Our method is not only competitive with other LLM-based and complex agentic frameworks but is also more scalable for real-world scenarios and on-device POI recommendations. Results demonstrate that our approach consistently outperforms baseline methods, offering a more interpretable and resource-efficient solution for POI recommendation systems. Our source code is available at: https://github.com/w11wo/GenUP.
Large language models, Location-based social networks, Point-of-interest recommendation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/YY/MM
1. Introduction
Location-based social networks (LBSNs) generate vast amounts of data, offering significant potential for modeling human mobility patterns. One application of LBSN data is next Point-of-Interest (POI) recommendation, where the goal is to predict a user’s likely next destination based on their historical check-ins. POI recommendation systems have evolved from early Markov models (Cheng et al., 2013; Rendle et al., 2010) to more advanced deep learning techniques, such as sequential models (Kong and Wu, 2018; Sun et al., 2020) and Graph Neural Networks (GNNs) (Lim et al., 2020; Zhang et al., 2022; Rao et al., 2022; Yang et al., 2022a). These models have significantly improved the ability to capture spatial-temporal dependencies and user mobility patterns. However, challenges remain, particularly with the cold-start problem, where limited interaction data makes it difficult to provide accurate predictions for inactive users. This is typically resolved by learning from similar historical trajectories from other users. More recently, large language models (LLMs) have been applied to next POI recommendation (Wang et al., 2023; Beneduce et al., 2024), leveraging their commonsense reasoning capabilities and their ability to incorporate contextual information. However, LLM-based methods can be computationally intensive due to their reliance on long historical trajectory prompts to capture cross-user patterns (Li et al., 2024). Similarly, agentic LLM frameworks (Feng et al., 2024), which employ multiple agents for different tasks, introduce further complexity and resource demands, making them difficult to implement in real-time applications.
Another challenge in POI recommendation is personalization, often achieved through dense vector representations of users. While these embeddings effectively model user behavior, they lack interpretability, transparency, and scrutability—qualities crucial for building user trust and improving recommendation explainability (O’Donovan and Smyth, 2005; Zhang and Chen, 2020). Natural language (NL) user profiles offer a promising alternative, where new users can quickly provide a concise description of their preferences, and existing users can update their profiles for more personalized and transparent recommendations (Radlinski et al., 2022). Despite their potential, NL profiles have not seen widespread adoption due to difficulties in evaluation and the ongoing development of methods for accurately generating these profiles (Ramos et al., 2024).
In response to these limitations, we propose a method that introduces NL user profiles into next POI recommendation systems. By replacing long historical trajectory prompts with concise NL profiles, our approach significantly reduces computational overhead and shortens the context length required for LLM-based recommendations. Furthermore, it improves personalization by shifting from dense user vector representations to interpretable NL profiles. We propose generating these profiles by analyzing LBSN check-ins to learn user routines, preferences, and behaviors. To create more grounded NL profiles, we integrate robust behavioral theories, capturing underlying intentions and patterns for more accurate POI predictions. In summary, our contributions are as follows:
  • We introduce natural language (NL) user profiles as system prompts to LLMs for next POI recommendation.
  • We leverage LBSN data to create NL user profiles, improving personalized recommendations and cross-user learning.
  • We propose a hybrid approach that combines the generation capabilities of LLMs with behavioral theories for more accurate user profiles.
2. Related Works
2.1. Large Language Models for Next POI Recommendation
Significant advancements in mobility prediction have been driven by deep learning models, moving from traditional Markov models (Cheng et al., 2013; Rendle et al., 2010) to more sophisticated approaches that capture high-order mobility patterns in trajectory data. Early works primarily relied on sequential-based methods like Recurrent Neural Networks (RNNs) and attention mechanisms to model individual user behavior. Models such as HST-LSTM (Kong and Wu, 2018) and LSTPM (Sun et al., 2020) were designed to capture spatio-temporal dependencies but lacked the ability to incorporate cross-user collaborative signals. This limitation prompted the development of graph-based models (Lim et al., 2020; Zhang et al., 2022; Rao et al., 2022; Yang et al., 2022a) to model relationships between users, locations, and trajectories. These models addressed the cold-start problem more effectively by using collaborative signals but struggled to integrate commonsense knowledge and broader contextual information (Li et al., 2024).
Large language models (LLMs) have since emerged as a viable solution for next POI recommendation, offering the ability to process both spatial-temporal data and commonsense knowledge. LLM-Mob (Wang et al., 2023), for example, introduced in-context learning and exploited the zero-shot reasoning capabilities of LLMs (Wei et al., 2022) to enhance next POI recommendation given the historical and contextual trajectories of a user. Similarly, LLM-ZS (Beneduce et al., 2024) simplified the zero-shot process and analyzed zero-shot, one-shot, and few-shot prompting scenarios. However, both methods depend heavily on task-specific prompting and require very large LMs, such as 70B instruction-tuned or closed-source LLMs, to achieve competitive performance. AgentMove (Feng et al., 2024) extends these techniques by employing multiple LLMs agents (Wang et al., 2024; Xi et al., 2023), each dedicated to different components of mobility prediction. This framework further integrates external knowledge sources such as urban structures, geospatial data, and location graphs, enabling the predictor LLM to process richer contextual information for the prediction process. While this multi-agent approach improves prediction accuracy and handles cold-start issues, it is computationally intensive and less suitable for real-time or resource-constrained environments.
LLM4POI (Li et al., 2024) proposed an alternative by conducting a supervised fine-tuning (SFT) of LLMs for next POI prediction, framing it as a question-answer task. This approach simplified the process by (1) using lighter 7B open-source LLMs, (2) streamlining in-context learning with trajectory prompting tailored for LBSN data, and (3) leveraging trajectory similarity from both current and other users to enhance the LLM’s predictive capabilities. While leveraging cross-user similar trajectories improves accuracy, LLM4POI’s reliance on long context lengths results in high computational costs. Building on this, our method introduces natural language user profiles to personalize POI recommendations and learn cross-user patterns from similar profiles, removing the need for long historical trajectory prompts and significantly decreasing the context length.
Figure 1. Our proposed framework of leveraging user check-in records as natural language user profiles as system prompts to large language models for supervised fine-tuning.
[…]
Our proposed framework of leveraging user check-in records as natural language user profiles as system prompts to large language models for supervised fine-tuning.
2.2. Natural Language User Profile for Recommender Systems
The use of natural language (NL) for user profiles in recommender systems marks a shift from traditional dense vector representations for user embeddings, addressing their limitations in transparency, interpretability, and scrutability (Radlinski et al., 2022). This shift is driven by the rapid adoption of large language models (LLMs), which excel at generating human-readable text, enabling the automated creation of NL user profiles.
Advantages of NL User Profiles
Radlinski et al. (2022) highlighted key advantages of NL user profiles over vector representations. NL user profiles (1) enhance transparency by allowing users to see their preferences, (2) improve interpretability by making preferences understandable, and (3) offer scrutability to modify their profiles.
NL profiles also help mitigate the cold-start problem, i.e. users who have not interacted much with a system. For example, new users can bootstrap their profiles by providing a concise textual description of their interests. Recommender systems can then leverage similar past profiles to infer similar patterns to generate recommendations (Radlinski et al., 2022). Furthermore, NL profiles enable users to dynamically steer recommendations as changes to their NL profile are immediately reflected in the recommendations they receive, providing greater control and trust (O’Donovan and Smyth, 2005) in the system (Ramos et al., 2024).
In terms of generalization, Ramos et al. (2024) and Zhou et al. (2024) have demonstrated that NL profiles are adaptable across domains including movie, TV show, and hotel reviews, proving effective on recommender tasks such as pairwise preference, pairwise choice, and rating prediction. This showcases the flexibility of NL in capturing a wide range of user preferences, from specific attributes like favorite genres, user expertise, and broader behavioral patterns.
Given these advantages, our method leverages LLMs to generate NL user profiles for point of interest (POI) prediction tasks. These profiles need to encompass not only the user’s basic information but also capture their routines, preferences, and behavioral patterns to enhance the accuracy of POI recommendations. Our primary user behavior data source will be user check-in records from location-based social networks (LBSNs). We hypothesize that this will uncover users’ patterns and routines, along with other relevant details to enhance the accuracy of POI recommendations. Unlike the other data sources mentioned, check-in records are unlabelled, exploiting large volumes of mobility data from online LBSNs that have been extensively collected and studied (Yang et al., 2015b, a, 2016; Cho et al., 2011; Feng et al., 2019).
Challenges in NL User Profile Generation
Despite the advantages, generating and optimizing NL profiles remains challenging. Profiles must be concise yet comprehensive, balancing short- and long-term preferences without overwhelming them with excessive information. Additionally, scaling NL profiles in large systems while avoiding biases in LLM-generated content remains difficult. Similarly, the evaluation of these profiles, in terms of fluency, coherence, and correctness, also lacks standardized metrics, complicating comparisons between methods (Radlinski et al., 2022; Zhou et al., 2024). (Ramos et al., 2024) for example, conducted a qualitative case study where the profiles were evaluated by humans. Further optimizations include reinforcement learning algorithms, such as REBEL (Gao et al., 2024) which iteratively refine NL profile generation by using recommendation quality metrics and employing a reward model.
Several approaches have emerged, from early template-based methods (Balog et al., 2019), to more advanced LLM-driven techniques (Brown et al., 2020). Hybrid methods, such as those combining template-based models with LLMs, have been employed to generate more nuanced and contextually relevant profiles, leveraging sentiment analysis and feature extraction (Ramos et al., 2024). Our method adopts a similar hybrid approach by grounding our NL profiles on user check-in records while also integrating robust personality assessment and behavioral theories such as the Theory of Planned Behavior (Ajzen, 1991; Shao et al., 2024), Big Five Inventory (BFI) traits (Briggs, 1992; De Raad, 2000; Goldberg, 2013; Jiang et al., 2023), and user attributes (Chen et al., 2024b). This combination enhances the accuracy of the generated NL profiles by capturing the user’s underlying desires, intentions, and behavioral patterns, which can lead to more accurate POI predictions.
3. Problem Definition
We adopted the same problem definition as LLM4POI (Li et al., 2024), converting sequences of user check-in records from LBSNs into prompts and treating next POI prediction as a supervised question-answering task. In their setup, each user check-in record is represented as a tuple $q=(u,p,c,t)$, where $u\in U$ with $U$ denoting the set of all users. The variable $p$ refers to a point of interest (POI) from the set $P=\left\{p_{1},p_{2},\dots,p_{M}\right\}$, where $M$ is the number of unique POIs. Additionally, $c$ denotes the POI’s category name, providing the LLM with extra semantic context about the POI, and $t$ is the timestamp of the check-in. We excluded the geometric coordinates (latitude, longitude) of the POIs, as they are irrelevant to the LLM (Li et al., 2024).
For each user $u$, trajectories were formed by grouping their check-in records into sequences within a specific time interval $\Delta t$. A trajectory $T_{i}^{u}(t)$ up to timestamp $t$ for user $u$ consists of a sequence of check-in:
$T_{i}^{u}(t)=\{(p_{1},c_{1},t_{1}),\dots,(p_{k},c_{k},t_{k})\},$
where $t_{1}<t_{2}<\dots<t_{k}=t$ and $t_{k}-t_{1}\leq\Delta t$. During training, the LLM is provided with a set of $L$ historical trajectories $\mathcal{T}_{u}=\{T_{1}^{u},T_{2}^{u},\dots,T_{L}^{u}\}$ for each user $u$, with the objective of predicting the next POI ID $p_{k+1}$, where the user will check in at the next timestamp $t_{k+1}$, given a current trajectory ${T^{\prime}}_{i}^{u}(t)$. The current trajectory refers to the list of POIs visited by the user prior to $t_{k+1}$ within the time interval $\Delta t$, while historical trajectories encompass previous trajectories occurring before the current time interval.
4. Methodology
Based on the findings of (Ramos et al., 2024) and building upon LLM4POI (Li et al., 2024), we hypothesize that NL user profiles can (1) replace long historical trajectories with a concise system prompt, (2) capture a user’s preferences and routines based on past check-ins, and (3) enable the LLM to learn from similar profiles to address the cold-start problem. Together, these capabilities aim to improve the efficiency and accuracy of next POI recommendation.
The overall framework we propose is illustrated in Figure 1. First, we generate NL user profiles from user check-in records, which are then used as system prompts for the LLM during the second stage of supervised fine-tuning (SFT). Unlike LLM4POI, our approach only used the user’s current trajectory as input, disregarding all historical trajectories from both the current and other users.
4.1. Natural Language User Profile Generation
We describe our proposed NL user profile generation framework as follows. We utilized large-scale, unsupervised user check-in records from location-based social networks (LBSNs), enriched with semantic information (POI category name), as our primary source of user behavior data. GPT-4o Mini [^1] We used gpt-4o-mini-2024-07-18 via the OpenAI API. was used as the large language model (LLM) to generate NL user profiles. To ground our NL user profile generation, we propose the integration of three components:
  • Theory of Planned Behaviour (Ajzen, 1991). As demonstrated by Shao et al. (2024), LLMs are capable of capturing a user’s attitude (preferences), subjective norms (routines), and perceived behavioral control (perceived likelihood of intentions based on profile, routine, and behavior history) from user activity logs/intention sequences. This improves the next activity and POI prediction. In our work, we focus on predicting user preferences and routines, such as frequently visited POIs and preferred POI categories.
  • Big Five Inventory (BFI) personality traits (Briggs, 1992; De Raad, 2000; Goldberg, 2013). Jiang et al. (2023) showed that pre-trained LLMs could generate content steered by the BFI personality traits they are asked to emulate, effectively emulating a specific user’s behavior and preferences. The BFI personality traits include: (1) extroverted/introverted, (2) agreeable/antagonistic, (3) conscientious/unconscientious, (4) neurotic/emotionally stable, and (5) open/closed to experience.
  • User Attributes. Chen et al. (2024b) showed that LLMs can reliably predict user attributes such as age, gender, educational background, and socioeconomic level. These attributes were intentionally selected as additional prompt information to the LLM as they are pivotal in shaping and making critical real-world decisions. Following TalkTuner (Chen et al., 2024b), we categorized the four attributes as follows: age (child: ¡13, adolescent: 13-17, adult: 18-64, older adult: ¿64); gender (male/female); educational background (some schooling, high school, college & beyond); and socioeconomic status (lower, middle, upper).
Following the approach of (Li et al., 2024) and using their prompt formatting structure, user check-in records and trajectories were first transformed into prompts. Trajectory prompting (Radford et al., 2021; Xue et al., 2022; Zhou et al., 2022) allows the conversion of heterogeneous data from LBSNs into natural language sentences that LLMs can understand. These trajectory prompts were then concatenated and used to generate NL user profiles, as outlined in Table 1. In addition to generating a 200-word NL user profile, the LLM was tasked with predicting the 3 components above. To facilitate easier parsing, the output is structured in JSON format. Examples of user profiles can be found in Appendix D.
Table 1. Natural language (NL) user profile generation prompt and check-in record trajectory prompt used for user profile generation. The teal text instructs the LLM to generate a 200-word user profile, orange for predicting preferences and routines, red for predicting BFI traits, and blue for predicting user attributes.
prompt Given the following check-ins of user [user id], generate a 200-word user profile summary to be used as a system prompt to another LLM that simulates this person’s behavior, preferences, routines, hobbies, schedule, etc. Predict this user’s Big Five Inventory traits: (1) extroverted / introverted, (2) agreeable / antagonistic, (3) conscientious / unconscientious, (4) neurotic / emotionally stable, (5) open / closed to experience. Also predict their age: child (¡13) / adolescent (13-17) / adult (18-64) / older adult (¿64); their gender: male / female; their educational background: some schooling / high school / college & beyond; and their socioeconomic level: lower / middle / upper. Also include any patterns that is observed and POI IDs of important places that might be visited in the future in the user profile summary. This system prompt will be used to make future check-in predictions. Return your response in JSON format: {”traits”: [trait1, trait2, trait3, trait4, trait5], ”attributes”: [age, gender, edu, socioeco], ”preferences”: [preference1, preference2, preference3, …], ”routines”: [routine1, routine2, routine3, …], ”user_profile”: ”200-word user profile summary”} [check-in records].
check-in record At [time], user [user id] visited POI id [poi id] which is a/an [poi category name] with category id [poi category id].
4.2. Trajectory Prompting with Natural Language User Profiles
Likewise, we followed a similar prompt structure to LLM4POI, with modifications to include NL user profiles and the use of Llama 2’s (Touvron et al., 2023b) chat prompt template, as shown in Table 2. Each prompt contains the user’s NL profile, the current trajectory block, an instruction block, and the corresponding target block. With LLM4POI’s setup, heterogenous LBSN data—such as user IDs, timestamps, and POI category names—are converted into sentences that capitalize on the LLM’s strong semantic understanding capabilities.
Table 2. Structure of NL user profile as system prompt, input prompt, and check-in record for supervised fine-tuning based on Llama 2’s chat prompt template. The teal text represents the system prompt, red for the current trajectory prompt, orange for the instruction prompt, and blue for the target block.
system prompt You are user [user id] and your basic information is as follows: Age: [age]; Gender: [gender]; Education: [education]; SocioEco: [socioeco]. You have the following traits: [trait1], [trait2], [trait3], [trait4], [trait5]. You have the following preferences: [preference1], […], [preferenceN]. You have the following routines: [routine1], […], [routineM]. [user profile]
prompt ¡s¿[INST] ¡¡SYS¿¿ [system prompt] ¡¡/SYS¿¿ The following is a trajectory of user [user id]: [check-in records]. Given the data, at time [time], which POI id will user [user id] visit? Note that POI id is an integer in the range from 0 to [POI id range]. [/INST] At [time], user [user id] will visit POI id [poi id]. ¡/s¿
check-in record At [time], user [user id] visited POI id [poi id] which is a/an [poi category name] with category id [poi category id].
The NL user profile generated using the method described in §4.1 serves as the system prompt for the LLM, providing a description of the user that the model is intended to simulate. Given that the output is formatted in JSON, with each component represented as a different key, it can be easily parsed and integrated into the input prompt. This system prompt informs the LLM of the user’s attributes, including age, gender, educational background, socioeconomic level, BFI traits, preferences, and routines. Consequently, the LLM’s POI prediction output can leverage the insights provided by the NL user profile. Additionally, the LLM can learn trajectory patterns and routines from users with similar NL profiles, thereby enhancing its predictive capabilities and helping to mitigate cold-start problems.
The current trajectory block consists of all check-in records from the current session, except for the latest record, which is placed in the target block. Unlike LLM4POI, we excluded historical check-in records from the user and other users to isolate the impact of our proposed NL user profiles and minimize the input context length. The range of POI IDs specified in the instruction block helps guide the LLM in generating a predicted POI ID, ensuring it falls within the set of POIs $P$. The target block, which includes the next timestamp, user ID, and next POI ID, serves as the generation targets/labels and is excluded from the input during the prediction process.
4.3. Supervised Fine-tuning
Moreover, we replicated a similar compute- and memory-efficient supervised fine-tuning (SFT) setup as LLM4POI (Li et al., 2024) to ensure a fair comparison with their approach as our baseline.
Parameter-Efficient Fine-tuning (PEFT)
Like LLM4POI, we applied LoRA (Hu et al., 2021) as the parameter-efficient fine-tuning (PEFT) technique used during training, which involves freezing all parameters in the pre-trained LLM and only training the newly injected low-rank decomposition matrices. These matrices are subsequently merged into the pre-trained LLM during the inference process. In their setup, LoRA matrices of rank 8 were injected only into the attention layers (query, key, value, and output projection layers) were injected with the new LoRA matrices of rank 8. This significantly reduces the GPU memory requirements when fine-tuning an LLM while maintaining similar performances to full fine-tuning.
Quantized Training
To further reduce GPU memory usage during the fine-tuning process, we applied the same quantization technique used by LLM4POI. Namely, we applied QLoRA (Quantized LoRA) (Dettmers et al., 2023) which combines LoRA with a novel 4-bit NormalFloat (NF4) data type for storing model weights, and applies double quantization. All forward and backward passes are still conducted in the BrainFloat16 (BF16) data type.
Efficient Training Kernels
As the trajectory prompts can be quite long after tokenization, we require a long context length of at least 16,384 tokens to fit our input prompt without truncation. LLM4POI utilized FlashAttention-2 (FA2) (Dao et al., 2022; Dao, 2023) in their training implementation to reduce the GPU memory usage and accommodate their 32,768-token sequence length. We similarly leveraged FA2 as our efficient scaled dot product attention (SDPA) kernel implementation, as well as the newly introduced Liger Kernel (Hsu et al., 2024) that increases training throughput and reduces memory usage. Like FA2, Liger Kernel introduced efficient GPU kernel implementations for other common LLM layers and operations such as RoPE (Su et al., 2023), RMSNorm (Zhang and Sennrich, 2019), SwiGLU (Shazeer, 2020), and cross-entropy loss.
5. Experiments
5.1. Setup
5.1.1. Datasets
As we aimed to compare our approach with that of LLM4POI, we used the same public datasets they used and employed their pre-processing code [^2] https://github.com/neolifer/LLM4POI. Specifically, we used the New York City (NYC) and Tokyo (TKY) Check-in Dataset (Yang et al., 2015b), which contains user check-ins from 12 April 2012 to 16 February 2013 from the popular LBSN, FourSquare [^3] https://foursquare.com/. They also used the Gowalla-CA dataset (Yan et al., 2023), a subset of the original Gowalla [^4] https://www.gowalla.com dataset (Cho et al., 2011), and filtered for check-in records from California and Nevada using geometric coordinates. Following their procedures, the filtered and pre-processed records are divided into training, validation, and test sets, using the first 80% for training, the middle 10% for validation, and the final 10% for testing. Only users and POIs that appear in the training set are kept in the validation and testing sets.
Furthermore, we recreated subsets of the Global-scale Check-in Dataset (Yang et al., 2019, 2022b), also derived from FourSquare, following the pre-processing methods outlined by (Feng et al., 2024). Trajectories were grouped into 72-hour sessions, with users having fewer than 5 sessions and sessions with fewer than 4 check-ins removed. The data was split into 70% for training, 10% for validation, and 20% for testing. Only users with between 3 and 50 sessions were included in the testing set. While they only used 200 test samples due to cost constraints, we evaluated on the entire testing set. We applied this pre-processing procedure to the cities of Moscow and São Paulo, two of the largest cities included in their evaluation. However, as shown in Appendix A, our reproduction of their dataset is not identical, as they filtered trajectories using an unspecified minimum distance from the city centers. In contrast, we used actual geometric boundaries of the cities obtained from OpenStreetMap [^5] https://www.openstreetmap.org and included only the POIs located within those boundaries. The actual geometric boundaries of cities are provided in the map located in Appendix B. All of the datasets we used are publicly available and anonymous.
5.1.2. Baselines
We compared our approach with the four kinds of baselines: supervised fine-tuned (SFT) LLMs, zero-shot in-context learning LLMs, agentic LLM frameworks, and graph-based models.
  • LLM4POI (Li et al., 2024): Main baseline, leveraging LLM for next POI recommendation by utilizing similar historical trajectories from current users and other users. Frames next POI prediction task as question-answer SFT pairs.
  • LLM-Mob (Wang et al., 2023): Employs in-context learning to enhance POI recommendations by providing historical and contextual user trajectories. Uses extensive prompting and zero-shot reasoning capabilities of LLMs.
  • LLM-ZS (Beneduce et al., 2024): A more simplified version of LLM-Mob, also utilizes historical and contextual trajectories, simpler prompting, and zero-shot reasoning. Compares zero-shot, one-shot, and few-shot prompting predictions.
  • AgentMove (Feng et al., 2024): Combines multiple LLMs as agents and integrates external knowledge sources such as urban structure, geospatial knowledge, and location graphs. Each agentic LLM serves as a different component, providing different information to the primary zero-shot, POI predictor LLM.
  • GETNext (Yang et al., 2022a): Graph-enhanced transformer model that incorporates a global trajectory flow map to leverage collaborative signals from other users and POI embeddings.
  • STHGCN (Yan et al., 2023): A spatio-temporal framework that models higher-order dependencies in time and space, using a hypergraph transformer to model cross-user relations.
5.1.3. Evaluation
We evaluated our model using Accuracy@1, the same metric employed by both LLM4POI and AgentMove, to assess the accuracy of predicting the next immediate POI.
5.1.4. Models
To ensure a fair comparison against the reported results of LLM4POI (Li et al., 2024), we used the same pre-trained LLM, Llama-2-7b-longlora-32k (Chen et al., 2024a), for our main comparative experiments. In their setup, the model without historical trajectories is denoted as LLM4POI*, while the model trained with historical trajectories from both the current user and other users is called LLM4POI. We denote our models with the prefix GenUP (generative user profilers).
Moreover, we fine-tuned more recent variants of the LLaMA family of models (Touvron et al., 2023a; Dubey et al., 2024), including Llama 3.1 8B and Llama 3.2 1B, to further benchmark our method against newer architectures and evaluate their performance on the next POI prediction task. The inclusion of the 1 billion model variant allows for a more efficient inference while maintaining similar performances and requiring limited computational resources, making it more accessible for researchers with constrained hardware.
5.1.5. Implementation
Our fine-tuning implementation details are provided in Appendix C, based on the setup from (Li et al., 2024). We used a linearly decaying learning rate scheduler for all fine-tuning runs, with 20 warmup steps and no weight decay. Since we excluded historical trajectories from our prompts, we used a context length of 16,384 tokens and trained for 3 epochs. The Transformer Reinforcement Learning (TRL) library (von Werra et al., 2020; Wolf et al., 2020) was used throughout our SFT experiments, which were conducted on NVIDIA H100 GPUs.
5.2. Results
We present our results in comparison to the baseline LLM4POI approach in Table 3. Additionally, we included their reported results for sequence-based and graph-based models to emphasize the performance of our proposed method against those that leverage additional historical trajectories of current users, as well as collaborative and intra-user social data.
Table 3. Evaluation results of our proposed method incorporating natural language user profiles as system prompts for LLM-based next POI prediction on New York City, Tokyo, and California datasets.
Model NYC TKY CA
Without historical and intra-user social data
LLM4POI* 0.2356 0.1517 0.1016
GenUP-Llama-2-7B 0.2575 0.1699 0.1094
GenUP-Llama-3.1-8B 0.2582 0.2127 0.1339
GenUP-Llama-3.2-1B 0.2484 0.1851 0.1267
With historical and intra-user social data
GETNext 0.2435 0.2254 0.1357
STHGCN 0.2734 0.2950 0.1730
LLM4POI 0.3372 0.3035 0.2065
The evaluation results from the NYC, TKY, and CA datasets demonstrate the effectiveness of our proposed method, which consistently outperforms LLM4POI*, both of which utilize the same Llama 2 base LLM. Specifically, GenUP-Llama-2-7B achieves higher accuracy in NYC (0.2575 vs. 0.2356), TKY (0.1699 vs. 0.1517), and CA (0.1094 vs. 0.1016). These findings indicate that incorporating natural language user profiles as system prompts improves performance, even when using the same pre-trained LLM.
Moreover, the larger GenUP-Llama-3.1-8B model, further increases performance, achieving the highest accuracy in all cities (NYC: 0.2582, TKY: 0.2127, CA: 0.1339). Notably, this model demonstrated competitive results compared to GETNext which exploits collaborative signals from cross-user data. Similarly, the smaller GenUP-Llama-3.2-1B model also consistently surpasses the LLM4POI* baseline, showing an improvement in accuracy despite using a smaller yet more powerful pre-trained base LLM.
Table 4. Evaluation results of our method compared to in-context learning and agentic LLM frameworks for next POI prediction on Moscow and São Paulo datasets. Results marked with are based on different dataset versions of the same city.
Model Base Model Moscow São Paulo
Supervised fine-tuning
LLM4POI* Llama 2 7B 0.146 0.166
GenUP-Llama-2-7B Llama 2 7B 0.159 0.175
GenUP-Llama-3.1-8B Llama 3.1 8B 0.163 0.178
GenUP-Llama-3.2-1B Llama 3.2 1B 0.161 0.175
In-context learning
LLM-Mob GPT-4o Mini 0.080 0.140
LLM-ZS GPT-4o Mini 0.120 0.165
Agentic LLM with external knowledge
AgentMove (base) GPT-4o Mini 0.120 0.165
AgentMove (+memory) GPT-4o Mini 0.160 0.190
AgentMove (+world) GPT-4o Mini 0.090 0.175
AgentMove (+collective) GPT-4o Mini 0.140 0.175
AgentMove (+mem+world) GPT-4o Mini 0.135 0.240
AgentMove (all) GPT-4o Mini 0.150 0.230
We also compare the results of our proposed method with the LLM-based agentic framework, such as AgentMove, which utilizes multiple LLMs as agents alongside external knowledge sources. Additionally, we include their reported results for other LLM-based baselines that rely on historical trajectories or employ in-context learning rather than SFT. We present our results in Table 4.
Although our dataset reproduction is not identical to that of (Feng et al., 2024), our models achieved competitive results on Moscow and São Paulo datasets compared to those that leverage in-context learning (ICL) methods, such as LLM-Mob and LLM-ZS. Both ICL approaches incorporate historical trajectories to inform their predictions, while AgentMove combines multiple LLM agents with external knowledge sources to enhance its predictive performance. In contrast, our method solely employs a single LLM base with supervised fine-tuning (SFT), highlighting the effectiveness of our approach even with a more straightforward architecture.
Notably, our models remain competitive with AgentMove, despite the latter (1) utilizing a much larger GPT-4o Mini model, (2) employing multiple LLM agents, and (3) incorporating external knowledge sources, such as a world knowledge generator and a collective knowledge extractor from cross-user mobility data. This further underscores the robustness of our approach, which achieves comparable performance with significantly fewer resources.
Likewise, our GenUP-Llama-2-7B model continues to outperform the LLM4POI* baseline across both datasets, while the GenUP-Llama-3.1-8B remains the top performer overall, even surpassing AgentMove with its extensive external knowledge components on the Moscow dataset. Additionally, the GenUP-Llama-3.2-1B model remains competitive with its larger counterparts, demonstrating comparable performance despite its reduced size.
5.3. Analysis
5.3.1. Impact of Natural Language User Profile Components
We investigated the impact of different components used to create our NL user profile, with the results shown in Table 5. In particular, we tested the following combination of components to include in the system prompt during training and inference:
  • No user profile: Baseline method of LLM4POI* with no system prompts (Li et al., 2024).
  • User Profile: A short, 200-word user profile.
  • User Profile + Routines & Preferences: The user profile along with the predicted routines and preferences based on the Theory of Planned Behavior (Ajzen, 1991).
  • User Profile + Routines & Preferences + Attributes: The user profile, predicted routines and preferences, and predicted user attributes (age, gender, educational background, socioeconomic status) (Chen et al., 2024b).
  • User Profile + Routines & Preferences + Attributes + BFI Traits: The full set, including the user profile, predicted routines and preferences, predicted user attributes, and predicted Big Five Inventory traits (Briggs, 1992; De Raad, 2000; Goldberg, 2013).
The results demonstrate that adding a simple, 200-word user profile generated from check-in records already improves the baseline performance of LLM4POI*, indicating that our approach of adding a system prompt enables the LLM to better emulate user behavior during POI prediction. However, not all predicted components lead to improved performance. For example, adding routines, preferences, and BFI traits does not impact the model’s predictive accuracy. On the other hand, the inclusion of predicted user attributes provides a slight improvement, suggesting that these attributes influence decision-making processes and user behavior, as proposed by (Chen et al., 2024b).
Table 5. Ablation study of the different natural language user profile components as system prompts on the NYC dataset.
System Prompt Model NYC
No user profile LLM4POI* 0.2356
200-word User Profile GenUP-Llama2-7B 0.2568
+ Routines & Preferences GenUP-Llama2-7B 0.2568
+ User Attributes GenUP-Llama2-7B 0.2575
+ BFI Traits GenUP-Llama2-7B 0.2575
5.3.2. User Cold-start Analysis
Inspired by (Li et al., 2024), we also conducted a user cold-start analysis on our proposed GenUP models across all five datasets in our study. Following their assessment method, users were categorized into different activity levels based on the number of trajectories in the training set. The top 30% of users were classified as active and the bottom 30% as inactive.
The results, presented in Table 6, illustrate a clear trend where the POI prediction accuracy increases with higher user activity levels, which aligns with intuitive expectations. For instance, the GenUP-Llama-2-7B model shows a progression in accuracy on the São Paulo dataset from 0.1366 for inactive users, to 0.1504 for normal users, and 0.1940 for very active users. This pattern holds across most models and cities, except for the CA dataset, where the models are slightly more accurate for inactive users, consistent with the findings reported by LLM4POI (Li et al., 2024).
Furthermore, larger models such as GenUP-Llama-3.1-8B consistently achieve better performance in very active user groups. For example, on the TKY dataset, GenUP-Llama3.1-8B achieves an accuracy of 0.2688, compared to 0.2063 for GenUP-Llama-2-7B. These results suggest that larger, more sophisticated models benefit from richer data available for very active users, leading to substantial gains in prediction accuracy. However, the variability in performance for inactive users indicates that the advantages of larger models are diminished when user activity data is limited.
Table 6. User cold-start analysis based on the number of trajectories on NYC, TKY, CA, Moscow and São Paulo datasets.
User Groups NYC TKY CA Moscow São Paulo
GenUP-Llama-2-7B
Inactive 0.2105 0.1306 0.1091 0.1227 0.1366
Normal 0.2591 0.1394 0.1089 0.1410 0.1504
Very Active 0.2752 0.2063 0.1096 0.1748 0.1940
GenUP-Llama-3.1-8B
Inactive 0.1826 0.1486 0.1380 0.1180 0.1393
Normal 0.2554 0.1695 0.1338 0.1464 0.1598
Very Active 0.2884 0.2688 0.1324 0.1808 0.1944
GenUP-Llama-3.2-1B
Inactive 0.1764 0.1306 0.1316 0.1210 0.1429
Normal 0.2664 0.1494 0.1223 0.1390 0.1530
Very Active 0.2704 0.2321 0.1263 0.1793 0.1906
5.3.3. Efficiency Comparison of Trajectory Similarity and User Profiling
Trajectory similarity computation is crucial for overcoming the cold-start problem, as it helps predict inactive users’ preferences through past historical trajectories. Traditional methods often use methods like Dynamic Time Warping (DTW), to measure point-to-point similarity between trajectories. For two trajectories with an average length of k, the time complexity of a single DTW comparison is $O(k^{2})$. When considering $L$ trajectories, the total complexity for comparing all pairs is $O(L^{2}\times k^{2})$. Additionally, sorting algorithms are typically required for ranking similarities, contributing an additional overhead of $O(L\log L)$, bringing the overall complexity to $O(L^{2}\times k^{2}+L\log L)$.
LLM4POI (Li et al., 2024) proposed to enhance predictive accuracy by leveraging historical trajectories but faces challenges, including a long context length of 32,768 tokens needed to fit the entire input prompt. Although their trajectory embedding approach reduces the complexity of trajectory embedding inference to $O(L)$, the necessity of performing pairwise similarity comparisons among $L$ trajectories results in a time complexity of $O(L^{2})$. With similarity sorting, the overall complexity becomes $O(L^{2}+L\log L)$. While this method improves prediction accuracy, it still incurs substantial computational costs, particularly in scenarios with frequent updates, such as when new trajectories are available and when user preferences shift.
In contrast, our method replaces frequent trajectory similarity computations by utilizing user profiles that require only a single inference per user given their historical trajectories $\mathcal{T}_{u}$. These profiles can be easily updated periodically as new trajectories become available, ensuring they reflect evolving user behaviors without constant recalibration. Consequently, the time complexity for generating these profiles is only $O(|U|)$ where $|U|$ is the number of users and $|U|\ll L$ This significant reduction in computational overhead, coupled with the ability to adapt efficiently to changes in user preferences, positions our method as a scalable solution for recommendation systems. Additionally, omitting historical trajectories from our prompt reduces the context length to 16,384 tokens, which is half of what LLM4POI required.
5.3.4. Open-source and Lightweight Base LLM
As shown in Table 7, our approach demonstrates competitive results compared to other LLM-based methods, all while utilizing open-source and lightweight models. Remarkably, the GenUP-Llama-3.2-1B model exhibited impressive predictive capabilities despite its smaller size, making it a promising candidate for offline or on-device POI recommendations, such as on mobile devices. Importantly, the reported results of LLM-ZS, LLM-Mob, and AgentMove in Table 4 used the closed-source GPT-4o Mini as their base LLM. Originally, LLM-Mob utilized GPT-3.5 Turbo (Wang et al., 2023), whereas LLM-ZS requires at least a 70 billion, instruction-tuned LLM or closed-source models like GPT-3.5, GPT-4, or GPT-4o to start exhibiting significant advantages (Beneduce et al., 2024). Although we also employed GPT-4o mini as our NL user profile generator in the experiments, this step can easily be replaced with open-source alternatives, such as instruction-tuned variants of the Llama 3.2 models that offer similar generation capabilities.
Table 7. Comparison of computational efficiency among our models and similar LLM-based methods, along with evaluation results on the NYC and TKY datasets. Results marked with $\sharp$, $\flat$, $\dagger$ are obtained from (Beneduce et al., 2024), (Wang et al., 2023), (Feng et al., 2024), respectively. Each of these studies utilized different versions of the same public dataset due to slight variations in their preprocessing steps.
Method Base Model #params NYC TKY Open
LLM-ZS Llama 2 Chat 7B 0.086 0.047
LLM-ZS Llama 2 7B 0.009 0.005
LLM-ZS Mistral 7B 0.087 0.069
LLM-ZS Llama 3 8B 0.011 0.009
LLM-ZS Llama 3 Instruct 8B 0.101 0.081
LLM-ZS Llama 3.1 8B 0.019 0.01
LLM-ZS Llama 2 Chat 13B 0.099 0.078
LLM-ZS Llama 2 13B 0.011 0.009
LLM-ZS Llama 2 Chat 70B 0.179 0.199
LLM-ZS Llama 2 70B 0.062 0.042
LLM-ZS Llama 3 70B 0.084 0.077
LLM-ZS Llama 3 Instruct 70B 0.196 0.199
LLM-Mob GPT-3.5 Turbo N/A 0.274 N/A $\times$
LLM-ZS GPT-3.5 Turbo N/A 0.192 0.199 $\times$
LLM-ZS GPT-4 N/A 0.195 0.205 $\times$
LLM-ZS GPT-4o N/A 0.199 0.209 $\times$
LLM-ZS GPT-4o Mini N/A N/A 0.175 $\times$
LLM-Mob GPT-4o Mini N/A N/A 0.175 $\times$
AgentMove GPT-4o Mini N/A N/A 0.185 $\times$
GenUP Llama 3.2 1B 0.248 0.185
GenUP Llama 2 7B 0.257 0.169
GenUP Llama 3.1 8B 0.258 0.212
6. Conclusion and Future Work
In this paper, we propose a novel approach for next point-of-interest (POI) recommendation by introducing natural language (NL) user profiles generated from location-based social network (LBSN) check-in records. Our method addresses key challenges in existing POI recommendation systems, such as cold-start issues and the lack of interpretability in user embeddings. By leveraging natural language user profiles, we aim to improve the personalization of recommendations, reduce computational complexity, and enhance transparency and user understanding of the recommendation process.
Our methodology consists of generating NL user profiles by analyzing users’ LBSN check-ins. We incorporate behavioral theories and user attributes to create a more grounded and comprehensive profile, capturing the user’s routines, preferences, and behavioral patterns. This approach eliminates the need for long historical trajectory prompts, allowing for more efficient context lengths in LLM-based recommendations. To validate our method, we conducted experiments on multiple LBSN datasets, comparing the performance of GenUP models with baseline methods, including models using long historical trajectories and agentic LLM frameworks. The results demonstrate that our approach outperforms traditional user embedding methods and agentic LLM frameworks, particularly in terms of personalization and efficiency. The analysis of these results highlights the benefits of using interpretable NL profiles, which provide a clearer understanding of user preferences and allow for easier adaptation and personalization in real-world applications. This makes our approach a promising solution for scalable, efficient, and transparent POI recommendation systems.
In the future, we plan to enhance NL user profiles by incorporating reinforcement learning to iteratively refine and optimize the generated profiles. Additionally, we aim to introduce intention modeling via reasoning, which would provide greater transparency and further improve the accuracy of recommendations by more accurately capturing users’ underlying motivations and decision-making processes.
Appendix A Dataset Statistics
We present the detailed statistics of our pre-processed version of the Global-scale Check-in Dataset in Table 8. AgentMove used an unspecified minimum distance to filter POI locations, whereas we used real geometric boundaries of cities to filter them.
Table 8. Dataset statistics comparing AgentMove’s version of next POI prediction datasets, derived from the Global-scale Check-in Dataset, with our reproduction. Differences arise from variations in POI location filtering procedures.
City Users Trajectories POIs Check-ins
AgentMove (Feng et al., 2024)
Moscow 10,501 100,854 93,599 950,898
São Paulo 11,856 77,120 78,904 809,198
Ours
Moscow 6,864 66,216 89,247 500,213
São Paulo 3,718 40,045 64,690 300,044
Appendix B Geometric Boundaries of Cities
We included maps that depict the geometric boundaries of Moscow and São Paulo used in our dataset filtering. Figure 2 presents the geometric boundaries of Moscow, while Figure 3 illustrates the boundaries of São Paulo.
Figure 2. Geometric boundaries of Moscow retrieved from OpenStreetMap.
[…]
Geometric boundaries of Moscow retrieved from OpenStreetMap.
Figure 3. Geometric boundaries of São Paulo retrieved from OpenStreetMap.
[…]
Geometric boundaries of São Paulo retrieved from OpenStreetMap.
Appendix C Hyperparameters
We show the different training hyperparameters of our LLMs across different datasets in Table 9. We also show the LoRA parameters we used for parameter-efficient fine-tuning in Table 10.
Table 9. Hyperparameters for GenUP models across different datasets. ”Others” represent the TKY, CA, Moscow, and São Paulo datasets.
Model Learning Batch Size
Rate NYC Others
GenUP-Llama-2-7B 2e-5 1 2
GenUP-Llama-3.1-8B 2e-4 1 2
GenUP-Llama-3.2-1B 2e-4 1 2
Table 10. LoRA parameters based on LLM4POI (Li et al., 2024).
parameter value
rank 8
alpha 16
dropout 0.0
bias none
modules query, key, value, output
Appendix D Sample User Profile Generation Results
We provide sample generation results for the user profiles from the NYC, TKY, and CA datasets in Figures 4, 5, and 6, respectively. We used GPT-4o Mini to generate these profiles given the check-in records of each user.
{minted}
[…]
Figure 4. Sample user profile of a user in the NYC dataset.
{minted}
[…]
Figure 5. Sample user profile of a user in the TKY dataset.
{minted}
[…]
Figure 6. Sample user profile of a user in the CA dataset.

### v2

@@
 Abstract.
-Traditional POI recommendation systems often lack transparency, interpretability, and scrutability due to their reliance on dense vector-based user embeddings.
+Traditional Point-of-Interest (POI) recommendation systems often lack transparency, interpretability, and scrutability due to their reliance on dense vector-based user embeddings.
 Furthermore, the cold-start problem—where systems have insufficient data for new users—limits their ability to generate accurate recommendations.
@@
 By incorporating NL profiles as system prompts to LLMs, our approach reduces reliance on extensive historical data, while remaining flexible, easily updated, and computationally efficient.
-Our method is not only competitive with other LLM-based and complex agentic frameworks but is also more scalable for real-world scenarios and on-device POI recommendations.
+Our method is not only competitive with other LLM-based and complex agentic frameworks but is also more scalable for real-world POI recommender systems.
 Results demonstrate that our approach consistently outperforms baseline methods, offering a more interpretable and resource-efficient solution for POI recommendation systems.
-Our source code is available at: https://github.com/w11wo/GenUP.
+Our source code is available at: https://github.com/w11wo/GenUP/.
 Large language models, Location-based social networks, Point-of-interest recommendation
@@
 One application of LBSN data is next Point-of-Interest (POI) recommendation, where the goal is to predict a user’s likely next destination based on their historical check-ins.
-POI recommendation systems have evolved from early Markov models (Cheng et al., 2013; Rendle et al., 2010) to more advanced deep learning techniques, such as sequential models (Kong and Wu, 2018; Sun et al., 2020) and Graph Neural Networks (GNNs) (Lim et al., 2020; Zhang et al., 2022; Rao et al., 2022; Yang et al., 2022a).
+POI recommendation systems have evolved from early Markov models (Cheng et al., 2013; Rendle et al., 2010) to more advanced deep learning techniques, such as sequential models (Kong and Wu, 2018; Sun et al., 2020) and Graph Neural Networks (Lim et al., 2020; Zhang et al., 2022; Rao et al., 2022; Yang et al., 2022a).
 These models have significantly improved the ability to capture spatial-temporal dependencies and user mobility patterns.
@@
 However, LLM-based methods can be computationally intensive due to their reliance on long historical trajectory prompts to capture cross-user patterns (Li et al., 2024).
-Similarly, agentic LLM frameworks (Feng et al., 2024), which employ multiple agents for different tasks, introduce further complexity and resource demands, making them difficult to implement in real-time applications.
+Similarly, agentic LLM frameworks (Feng et al., 2024), which employ multiple agents for different tasks introduce added complexity and resource demands.
 Another challenge in POI recommendation is personalization, often achieved through dense vector representations of users.
@@
 Despite their potential, NL profiles have not seen widespread adoption due to difficulties in evaluation and the ongoing development of methods for accurately generating these profiles (Ramos et al., 2024).
-In response to these limitations, we propose a method that introduces NL user profiles into next POI recommendation systems.
-By replacing long historical trajectory prompts with concise NL profiles, our approach significantly reduces computational overhead and shortens the context length required for LLM-based recommendations.
-Furthermore, it improves personalization by shifting from dense user vector representations to interpretable NL profiles.
-We propose generating these profiles by analyzing LBSN check-ins to learn user routines, preferences, and behaviors.
-To create more grounded NL profiles, we integrate robust behavioral theories, capturing underlying intentions and patterns for more accurate POI predictions.
-In summary, our contributions are as follows:
-  • We introduce natural language (NL) user profiles as system prompts to LLMs for next POI recommendation.
-  • We leverage LBSN data to create NL user profiles, improving personalized recommendations and cross-user learning.
-  • We propose a hybrid approach that combines the generation capabilities of LLMs with behavioral theories for more accurate user profiles.
+Prior work, such as LLMob (Wang et al., 2024a), relies on rigid user categorization, generating multiple candidate personas and selecting the most likely one based on a given trajectory.
+While effective, this approach limits personalization by constraining users to predefined categories.
+In response, we introduce NL user profiles for next POI recommendation, replacing long historical trajectory prompts with concise profiles to reduce computational overhead and context length.
+We propose generating these profiles from LBSN check-ins to learn user behaviors and integrate robust behavioral theories for more accurate POI predictions.
+Our contributions are as follows:
+  • We introduce natural language (NL) user profiles as LLM system prompts for next POI recommendation.
+  • We reconstruct NL user profiles from LBSN data, capturing individualized routines without predefined categories.
+  • We combine LLM generation and behavioral theories for more grounded user profiles.
+Our method enhances POI recommendation by improving transparency, reducing reliance on extensive historical data, and maintaining competitive accuracy with a more scalable, resource-efficient design.
+Unlike existing LLM-based approaches, it achieves strong performance while using smaller, open-source models, making it more adaptable for real-world deployment.
 2. Related Works
@@
 However, both methods depend heavily on task-specific prompting and require very large LMs, such as 70B instruction-tuned or closed-source LLMs, to achieve competitive performance.
-AgentMove (Feng et al., 2024) extends these techniques by employing multiple LLMs agents (Wang et al., 2024; Xi et al., 2023), each dedicated to different components of mobility prediction.
+AgentMove (Feng et al., 2024) extends these techniques by employing multiple LLMs agents (Wang et al., 2024b; Xi et al., 2023), each dedicated to different components of mobility prediction.
 This framework further integrates external knowledge sources such as urban structures, geospatial data, and location graphs, enabling the predictor LLM to process richer contextual information for the prediction process.
@@
 4.1. Natural Language User Profile Generation
+Generation
 We describe our proposed NL user profile generation framework as follows.
@@
 To facilitate easier parsing, the output is structured in JSON format.
-Examples of user profiles can be found in Appendix D.
+Examples of a user profile can be found in Appendix A.
+Validation and Challenges
+A key challenge in user profile generation is validation, as extensively discussed in §2.2.
+Another challenge we encountered was LLM hallucinations, where it failed to follow the instructed JSON schema, leading to inconsistent parsing due to changed JSON key names.
+However, this issue affected fewer than five users and was easily resolved by re-running the prompt.
+To validate our generated profiles, we ensured that predicted personality traits were valid BFI categories and that user attributes matched predefined classes.
+Additionally, we employed the LLM-as-a-Judge method (Zheng et al., 2023), where LLMs assessed the quality of generated profiles.
+While this approach has potential biases, we consider it a reasonable proxy for validation and have provided further details in Appendix B.
+Interestingly, our user profiles are also weak in conciseness, parallel with the findings of (Ramos et al., 2024), who employed human evaluators.
+To further evaluate the biases in our generated NL user profiles, we present their distribution in Appendix C.
 Table 1.
@@
 Consequently, the LLM’s POI prediction output can leverage the insights provided by the NL user profile.
-Additionally, the LLM can learn trajectory patterns and routines from users with similar NL profiles, thereby enhancing its predictive capabilities and helping to mitigate cold-start problems.
+Additionally, the LLM can learn trajectory patterns and routines from users with similar NL profiles, thereby enhancing its predictive capabilities.
 The current trajectory block consists of all check-in records from the current session, except for the latest record, which is placed in the target block.
 Unlike LLM4POI, we excluded historical check-in records from the user and other users to isolate the impact of our proposed NL user profiles and minimize the input context length.
+[^2] Our method only implicitly incorporates historical information into its user profile, unlike LLM4POI.
+The latter has access to multiple similar, repeatable trajectories, whereas the former can only summarize frequent or long-term trajectories into a concise user profile.
+LLM4POI also incorporates intra-user social data, while ours does not.
 The range of POI IDs specified in the instruction block helps guide the LLM in generating a predicted POI ID, ensuring it falls within the set of POIs $P$.
@@
 These matrices are subsequently merged into the pre-trained LLM during the inference process.
-In their setup, LoRA matrices of rank 8 were injected only into the attention layers (query, key, value, and output projection layers) were injected with the new LoRA matrices of rank 8.
+In their setup, LoRA matrices of rank 8 were injected into the attention layers (query, key, value, and output projection layers).
 This significantly reduces the GPU memory requirements when fine-tuning an LLM while maintaining similar performances to full fine-tuning.
@@
 We applied this pre-processing procedure to the cities of Moscow and São Paulo, two of the largest cities included in their evaluation.
-However, as shown in Appendix A, our reproduction of their dataset is not identical, as they filtered trajectories using an unspecified minimum distance from the city centers.
+However, as shown in Appendix D, our reproduction of their dataset is not identical, as they filtered trajectories using an unspecified minimum distance from the city centers.
 In contrast, we used actual geometric boundaries of the cities obtained from OpenStreetMap [^6] https://www.openstreetmap.org and included only the POIs located within those boundaries.
-The actual geometric boundaries of cities are provided in the map located in Appendix B.
 All of the datasets we used are publicly available and anonymous.
@@
 Uses extensive prompting and zero-shot reasoning capabilities of LLMs.
-  • LLM-ZS (Beneduce et al., 2024): A more simplified version of LLM-Mob, also utilizes historical and contextual trajectories, simpler prompting, and zero-shot reasoning.
+  • LLM-ZS (Beneduce et al., 2024): A simplified version of LLM-Mob, also utilizes historical and contextual trajectories, simpler prompts, and zero-shot reasoning.
 Compares zero-shot, one-shot, and few-shot prompting predictions.
@@
 5.1.3. Evaluation
-We evaluated our model using Accuracy@1, the same metric employed by both LLM4POI and AgentMove, to assess the accuracy of predicting the next immediate POI.
+We evaluated our model using Accuracy@1, the same metric employed by both LLM4POI and AgentMove, to assess prediction accuracy when a user can only visit one next POI rather than multiple locations.
 5.1.4. Models
@@
 5.1.5. Implementation
-Our fine-tuning implementation details are provided in Appendix C, based on the setup from (Li et al., 2024).
+Our fine-tuning implementation details are provided in Appendix E, based on the setup from (Li et al., 2024).
 We used a linearly decaying learning rate scheduler for all fine-tuning runs, with 20 warmup steps and no weight decay.
@@
 Additionally, we included their reported results for sequence-based and graph-based models to emphasize the performance of our proposed method against those that leverage additional historical trajectories of current users, as well as collaborative and intra-user social data.
-Table 3. Evaluation results of our proposed method incorporating natural language user profiles as system prompts for LLM-based next POI prediction on New York City, Tokyo, and California datasets.
+Table 3.
+Evaluation results of our proposed method incorporating natural language user profiles as system prompts for LLM-based next POI prediction on New York City, Tokyo, and California datasets.
+[^8] LM4POI* refers to the baseline method without historical trajectories (§5.1.4).
+For clarity, the highest score on each dataset is highlighted in bold, while the second highest is underlined.
 Model NYC TKY CA
@@
 These findings indicate that incorporating natural language user profiles as system prompts improves performance, even when using the same pre-trained LLM.
-Moreover, the larger GenUP-Llama-3.1-8B model, further increases performance, achieving the highest accuracy in all cities (NYC: 0.2582, TKY: 0.2127, CA: 0.1339).
+Moreover, the larger GenUP-Llama-3.1-8B model, further increases performance, achieving the highest accuracy in all three cities (NYC: 0.2582, TKY: 0.2127, CA: 0.1339).
 Notably, this model demonstrated competitive results compared to GETNext which exploits collaborative signals from cross-user data.
@@
 Evaluation results of our method compared to in-context learning and agentic LLM frameworks for next POI prediction on Moscow and São Paulo datasets.
-Results marked with are based on different dataset versions of the same city.
+Results marked with are based on different dataset versions of the same city.footnotemark:
 Model Base Model Moscow São Paulo
@@
 On the other hand, the inclusion of predicted user attributes provides a slight improvement, suggesting that these attributes influence decision-making processes and user behavior, as proposed by (Chen et al., 2024b).
-Table 5. Ablation study of the different natural language user profile components as system prompts on the NYC dataset.
+Table 5. Ablation study of the different natural language user profile components as system prompts on the NYC dataset.footnotemark:
 System Prompt Model NYC
@@
 However, the variability in performance for inactive users indicates that the advantages of larger models are diminished when user activity data is limited.
+We also analyzed how different NL user profile components affect different user groups, with results shown in Table 7.
+Our findings indicate that their impact varies by user activity level.
+For inactive (cold-start) users, incorporating routines and preferences improved accuracy by +0.93%, with user attributes providing another +0.93% boost.
+However, while adding BFI traits reduced accuracy for inactive users – as also observed in §5.3.1 – they enhanced the performance for highly active users.
 Table 6. User cold-start analysis based on the number of trajectories on NYC, TKY, CA, Moscow and São Paulo datasets.
@@
 Very Active 0.2704 0.2321 0.1263 0.1793 0.1906
+Table 7. Impact of different NL user profile components on next POI recommendation accuracy for NYC users, categorized by activity level, using the GenUP-Llama2-7B model.
+System Prompt Inactive Normal Very Active
+200-word User Profile 0.2012 0.2737 0.2728
++ Routines & Preferences 0.2105 0.2774 0.2680
++ User Attributes 0.2198 0.2664 0.2692
++ BFI Traits 0.2105 0.2591 0.2752
 5.3.3. Efficiency Comparison of Trajectory Similarity and User Profiling
@@
 In contrast, our method replaces frequent trajectory similarity computations by utilizing user profiles that require only a single inference per user given their historical trajectories $\mathcal{T}_{u}$.
+Consequently, the time complexity for generating these profiles is only $O(|U|)$ where $|U|$ is the number of users and $|U|\ll L$.
 These profiles can be easily updated periodically as new trajectories become available, ensuring they reflect evolving user behaviors without constant recalibration.
-Consequently, the time complexity for generating these profiles is only $O(|U|)$ where $|U|$ is the number of users and $|U|\ll L$ This significant reduction in computational overhead, coupled with the ability to adapt efficiently to changes in user preferences, positions our method as a scalable solution for recommendation systems.
-Additionally, omitting historical trajectories from our prompt reduces the context length to 16,384 tokens, which is half of what LLM4POI required.
+An example is to employ a user profile embedding shift, where a new profile $P^{\prime}$ replaces an existing profile $P$ if their similarity falls below a predefined threshold, indicating a significant behavioral shift.
+This strategy prioritizes meaningful long-term changes over fixed update intervals while allowing short-term adaptations when necessary.
+In practice, user profile generation with GPT-4o Mini averaged under 2 seconds per user and can be further accelerated with batched and parallel inference, making updates efficient even at scale.
+Notably, our method omits historical trajectories halving the context length of LLM4POI.
+This significantly reduces memory and compute requirements.
 5.3.4. Open-source and Lightweight Base LLM
-As shown in Table 7, our approach demonstrates competitive results compared to other LLM-based methods, all while utilizing open-source and lightweight models.
+As shown in Appendix F, our approach demonstrates competitive results compared to other LLM-based methods, all while utilizing open-source and lightweight models.
 Remarkably, the GenUP-Llama-3.2-1B model exhibited impressive predictive capabilities despite its smaller size, making it a promising candidate for offline or on-device POI recommendations, such as on mobile devices.
@@
 Although we also employed GPT-4o mini as our NL user profile generator in the experiments, this step can easily be replaced with open-source alternatives, such as instruction-tuned variants of the Llama 3.2 models that offer similar generation capabilities.
+6. Conclusion and Future Work
+In this paper, we propose a novel approach for next point-of-interest (POI) recommendation by introducing natural language (NL) user profiles generated from location-based social network (LBSN) check-in records.
+Our method addresses key challenges in existing POI recommendation systems, such as cold-start issues and the lack of interpretability in user embeddings.
+By leveraging natural language user profiles, we aim to improve the personalization of recommendations, reduce computational complexity, and enhance transparency and user understanding of the recommendation process.
+Our methodology consists of generating NL user profiles by analyzing users’ LBSN check-ins.
+We incorporate behavioral theories and user attributes to create a more grounded and comprehensive profile, capturing the user’s routines, preferences, and behavioral patterns.
+This approach eliminates the need for long historical trajectory prompts, allowing for more efficient context lengths in LLM-based recommendations.
+To validate our method, we conducted experiments on multiple LBSN datasets, comparing the performance of GenUP models with baseline methods, including models using long historical trajectories and agentic LLM frameworks.
+The results demonstrate that our approach outperforms traditional user embedding methods and agentic LLM frameworks, particularly in terms of personalization and efficiency.
+The analysis of these results highlights the benefits of using interpretable NL profiles, which provide a clearer understanding of user preferences and allow for easier adaptation and personalization in real-world applications.
+This makes our approach a promising solution for scalable, efficient, and transparent POI recommendation systems.
+Acknowledgements.
+We would like to thank the support of the National Computational Infrastructure and the ARC Center of Excellence for Automated Decision Making and Society (CE200100005).
+Computational facilities were provided by the School of Computer Science and Engineering at UNSW Sydney through the Wolfpack computational cluster.
+We further thank OpenAI’s Researcher Access Program for API access to GPT models and the Google Cloud Research Credits program for their support.
+Appendix A Sample User Profile Generation Results
+We provide a sample generation result of a user profile from the NYC dataset in Figure 2. We used GPT-4o Mini to generate profiles given the check-in records of each user.
+Appendix B User Profile Validation
+We followed (Ramos et al., 2024) and their qualitative case study on NL profiles, adopting their metrics to assess our generated user profiles.
+LLMs were prompted using the validation prompt shown in Table 8, with evaluation results presented in Table 9.
+Table 8. User profile validation prompt provided to an LLM for evaluation, using metrics from (Ramos et al., 2024).
+prompt Given the following natural language (NL) user profile [user profile], you are to assess the user profile based on four criteria:
+- Fluency: Is the NL profile both syntactically and semantically correct?
+- Informativeness: Does the NL profile provide important information for a user profile?
+- Conciseness: Is the NL profile written in a concise manner?
+- Relevance: Given the list of reviews, is the NL profile relevant to the user?
+Return your response in the JSON format: {”fluency”: 0/1, ”informativeness”: 0/1, ”conciseness”: 0/1, ”relevance”: 0/1}
+Table 9. Evaluation results of LLM-judged user profile validation using metrics from (Ramos et al., 2024).
+Metric NYC CA TKY
+Judge: GPT 4-o Mini
+Fluency 100.00 100.00 100.00
+Informativeness 100.00 100.00 100.00
+Conciseness 43.17 37.26 40.60
+Relevance 100.00 100.00 100.00
+Judge: Gemini 1.5 Flash
+Fluency 100.00 100.00 100.00
+Informativeness 100.00 100.00 100.00
+Conciseness 65.81 57.68 63.48
+Relevance 100.00 100.00 100.00
+Appendix C User Profile Distribution
+We present the distribution of the generated NL user profiles across three datasets: NYC, CA, and TKY.
+Table 10 displays the distribution of predicted personality traits, while Table 11 shows the distribution of user attributes, highlighting the diversity within each dataset.
+Table 10. Distribution of predicted BFI traits across the NYC, CA, and TKY datasets.
+NYC CA TKY
+Extroverted 736 3526 1476
+Introverted 311 430 805
+Agreeable 1047 3956 2281
+Antagonistic 0 0 0
+Conscientious 1047 3956 2281
+Unconscientious 0 0 0
+Neurotic 41 268 99
+Emotionally Stable 1006 3688 2182
+Open to Experience 1043 3947 2279
+Closed to Experience 4 9 2
+Table 11. Distribution of predicted user attributes across the NYC, CA, and TKY datasets.
+NYC CA TKY
+Age Child 1 14 2
+Adolescent 72 330 274
+Adult 974 3612 2005
+Older Adult 0 0 0
+Gender Male 774 2911 1881
+Female 273 1045 400
+Education Some Schooling 43 190 192
+High School 55 176 165
+College & Beyond 949 3590 1924
+Socioeconomic Lower 3 7 3
+Middle 1028 3651 2240
+Upper 16 298 38
+Appendix D Dataset Statistics
+We present the detailed statistics of our pre-processed version of the Global-scale Check-in Dataset in Table 12.
+AgentMove used an unspecified minimum distance to filter POI locations, whereas we used real geometric boundaries of cities to filter them.
+Table 12.
+Dataset statistics comparing AgentMove’s version of next POI prediction datasets, derived from the Global-scale Check-in Dataset, with our reproduction.
+Differences arise from variations in POI location filtering procedures.
+City Users Trajectories POIs Check-ins
+AgentMove (Feng et al., 2024)
+Moscow 10,501 100,854 93,599 950,898
+São Paulo 11,856 77,120 78,904 809,198
+Ours
+Moscow 6,864 66,216 89,247 500,213
+São Paulo 3,718 40,045 64,690 300,044
+Appendix E Hyperparameters
+We show the different training hyperparameters of our LLMs across different datasets in Table 13. We also show the LoRA parameters we used for parameter-efficient fine-tuning in Table 14.
+Table 13. Hyperparameters for GenUP models across different datasets. ”Others” represent the TKY, CA, Moscow, and São Paulo datasets.
+Model Learning Batch Size
+Rate NYC Others
+GenUP-Llama-2-7B 2e-5 1 2
+GenUP-Llama-3.1-8B 2e-4 1 2
+GenUP-Llama-3.2-1B 2e-4 1 2
+Table 14. LoRA parameters based on LLM4POI (Li et al., 2024).
+parameter value
+rank 8
+alpha 16
+dropout 0.0
+bias none
+modules query, key, value, output
+Appendix F Computational Efficiency Comparison
+In Table 15, we compare the base LLMs used in our approach with those employed by existing methods.
+Our method leverages open-source, lightweight models, while previous works such as LLM-ZS (Beneduce et al., 2024), LLM-Mob (Wang et al., 2023), and AgentMove (Feng et al., 2024) rely on larger and/or or closed-source LLMs.
 Table 15.
@@
 GenUP Llama 3.1 8B 0.258 0.212
-6. Conclusion and Future Work
-In this paper, we propose a novel approach for next point-of-interest (POI) recommendation by introducing natural language (NL) user profiles generated from location-based social network (LBSN) check-in records.
-Our method addresses key challenges in existing POI recommendation systems, such as cold-start issues and the lack of interpretability in user embeddings.
-By leveraging natural language user profiles, we aim to improve the personalization of recommendations, reduce computational complexity, and enhance transparency and user understanding of the recommendation process.
-Our methodology consists of generating NL user profiles by analyzing users’ LBSN check-ins.
-We incorporate behavioral theories and user attributes to create a more grounded and comprehensive profile, capturing the user’s routines, preferences, and behavioral patterns.
-This approach eliminates the need for long historical trajectory prompts, allowing for more efficient context lengths in LLM-based recommendations.
-To validate our method, we conducted experiments on multiple LBSN datasets, comparing the performance of GenUP models with baseline methods, including models using long historical trajectories and agentic LLM frameworks.
-The results demonstrate that our approach outperforms traditional user embedding methods and agentic LLM frameworks, particularly in terms of personalization and efficiency.
-The analysis of these results highlights the benefits of using interpretable NL profiles, which provide a clearer understanding of user preferences and allow for easier adaptation and personalization in real-world applications.
-This makes our approach a promising solution for scalable, efficient, and transparent POI recommendation systems.
-In the future, we plan to enhance NL user profiles by incorporating reinforcement learning to iteratively refine and optimize the generated profiles.
-Additionally, we aim to introduce intention modeling via reasoning, which would provide greater transparency and further improve the accuracy of recommendations by more accurately capturing users’ underlying motivations and decision-making processes.
-Appendix A Dataset Statistics
-We present the detailed statistics of our pre-processed version of the Global-scale Check-in Dataset in Table 8.
-AgentMove used an unspecified minimum distance to filter POI locations, whereas we used real geometric boundaries of cities to filter them.
-Table 8.
-Dataset statistics comparing AgentMove’s version of next POI prediction datasets, derived from the Global-scale Check-in Dataset, with our reproduction.
-Differences arise from variations in POI location filtering procedures.
-City Users Trajectories POIs Check-ins
-AgentMove (Feng et al., 2024)
-Moscow 10,501 100,854 93,599 950,898
-São Paulo 11,856 77,120 78,904 809,198
-Ours
-Moscow 6,864 66,216 89,247 500,213
-São Paulo 3,718 40,045 64,690 300,044
-Appendix B Geometric Boundaries of Cities
-We included maps that depict the geometric boundaries of Moscow and São Paulo used in our dataset filtering.
-Figure 2 presents the geometric boundaries of Moscow, while Figure 3 illustrates the boundaries of São Paulo.
-Figure 2. Geometric boundaries of Moscow retrieved from OpenStreetMap.
-[…]
-Geometric boundaries of Moscow retrieved from OpenStreetMap.
-Figure 3. Geometric boundaries of São Paulo retrieved from OpenStreetMap.
-[…]
-Geometric boundaries of São Paulo retrieved from OpenStreetMap.
-Appendix C Hyperparameters
-We show the different training hyperparameters of our LLMs across different datasets in Table 9. We also show the LoRA parameters we used for parameter-efficient fine-tuning in Table 10.
-Table 9. Hyperparameters for GenUP models across different datasets. ”Others” represent the TKY, CA, Moscow, and São Paulo datasets.
-Model Learning Batch Size
-Rate NYC Others
-GenUP-Llama-2-7B 2e-5 1 2
-GenUP-Llama-3.1-8B 2e-4 1 2
-GenUP-Llama-3.2-1B 2e-4 1 2
-Table 10. LoRA parameters based on LLM4POI (Li et al., 2024).
-parameter value
-rank 8
-alpha 16
-dropout 0.0
-bias none
-modules query, key, value, output
-Appendix D Sample User Profile Generation Results
-We provide sample generation results for the user profiles from the NYC, TKY, and CA datasets in Figures 4, 5, and 6, respectively.
-We used GPT-4o Mini to generate these profiles given the check-in records of each user.
 {minted}
@@
 Figure 2. Sample user profile of a user in the NYC dataset.
-{minted}
-[…]
-Figure 5. Sample user profile of a user in the TKY dataset.
-{minted}
-[…]
-Figure 6. Sample user profile of a user in the CA dataset.